# ATLAS: Automated Toolkit for Large-Scale Verified Code Synthesis

### v1

@@
Abstract.
Large language models have shown potential for program verification, but progress is hindered by the scarcity of verified code for training. We present ATLAS [^1] Automated Toolkit for LArge-Scale Verified Code Synthesis, an automated pipeline that synthesizes verified programs at scale to address this data bottleneck. ATLAS generates complete Dafny programs with specifications, implementations, and proofs, producing 2.7K verified programs from which we extract over 19K training examples—more than 7 per verified program—by decomposing the synthesis process into multiple specialized tasks. Fine-tuning Qwen 2.5 7B Coder on this dataset produces substantial gains: +23 percentage points on DafnyBench and +50 percentage points on DafnySynthesis. These results demonstrate that synthetic verified code can effectively enhance LLM capabilities for formal verification.
copyright: noneconference: ; 11th January, 2026; Rennes, France
1. Introduction
Figure 1. An overview of the ATLAS pipeline.
Large language models (LLMs) have had a significant impact on the landscape of software engineering, demonstrating sophisticated capabilities in code generation, completion, and debugging across a variety of programming languages. This impact has been most notable on mainstream languages such as Python, JavaScript, and Java (chen2021codex; cassano2023multiple; kocetkov2022stack) for which training data is abundant. This success has sparked significant interest in the use of these models in the more challenging domain of formal verification—the process of mathematically proving the correctness of software. Currently, each verified program represents a substantial time investment of human experts to craft not only the implementation but also the intricate specifications, assertions, and invariants that proof assistants require. However, what if generative systems could assist human experts—suggesting specifications, automating proof steps, and synthesizing verified code—to make verified software more attainable?
Unfortunately, the effective application of LLMs to auto-active verification languages like Dafny (dafny), SPARK (SPARK), F (fstar), and Viper (viper) is constrained by a significant data scarcity problem. Although some source code is openly available on, for example, GitHub, formally verified code remains rare compared to code in general-purpose languages. Fortunately, the challenge is not unique to auto-active verification languages, and the formal mathematics methods research community has seen promising results in overcoming similar data hurdles for interactive theorem provers. Recent work (polu2022formalmathematicsstatementcurriculum; ying2025leanworkbooklargescalelean; wang2025kiminaproverpreviewlargeformal; lin2025goedelproverfrontiermodelopensource) in the Lean ecosystem, for instance, has demonstrated that machine-generated data, in the form of autoformalized theorem statements, can be successfully used to train models for complex theorem-proving tasks. Inspired by these successes, we set out to address the data scarcity problem for Dafny.
In this paper, we introduce ATLAS, a novel pipeline designed to automatically generate a large dataset of Dafny programs that verify against non-trivial specifications (Figure 1). By executing our pipeline on the TACO-verified dataset (li2023tacotopicsalgorithmiccode), which consists of LeetCode-style problems with Python solutions and unit tests, we were able to generate 2.7K new high-quality verified Dafny programs, and consequently extract around 19K data points for model training (Table 1). To validate our approach, we fine-tuned the open-source Qwen 2.5 7B Coder model on our synthesized dataset, observing strong improvements on the popular DafnyBench (loughridge2024dafnybenchbenchmarkformalsoftware) and DafnySynthesis (misu2024towards) benchmarks (Figure 2).
@@ 1. Introduction
  • We present ATLAS, an automated pipeline for synthesizing verified Dafny programs at scale.
  • We show empirically that fine-tuning on our synthesized data significantly improves LLM performance on formal verification tasks in Dafny.
  • We define soundness and completeness criteria and integrate them with LLM-as-a-Judge feedback to refine both specifications and implementations.
  • We maximize data efficiency by decomposing the output of ATLAS into specialized training tasks, generating over $7$ training examples per verified program on average.
Figure 2. Evaluation of our fine-tuned Qwen 2.5 7B Coder on DafnyBench and DafnySynthesis, compared with results reported in (loughridge2024dafnybenchbenchmarkformalsoftware) and (misu2024towards) at the time of their release.
2. ATLAS
@@
To address these requirements, we designed ATLAS, a multi-phase pipeline that decouples problem statement generation from verified solution synthesis. The complete process is shown in Figure 1.
2.1. Overview
@@
Figure 3. A verified Dafny program synthesized by ATLAS, including specification, implementation, and test cases.
We seed the synthesis pipeline with a dataset of programming contest problems paired with solutions in a general-purpose language. This ensures each program has a self-contained implementation and reference solution with input-output test cases. We use TACO-verified (li2023tacotopicsalgorithmiccode), which contains 12.8K programming contest problems with Python reference solutions. We selected TACO-verified for its breadth—covering problems from LeetCode, CodeForces, HackerRank, and others—its unified and rich metadata, and its permissive licensing. Other datasets that could potentially be used for this task include APPS (APPS) and CodeContests (CodeConstests), which offer similar problem-solution pairs. However, they may lack the same high-quality metadata and strict test-case quality that TACO-verified provides.
2.1.1. Contract Generation
@@
The Dafny compiler verifies generated specifications and tests for syntax errors. On failure, compiler errors are formatted and passed to an LLM judge, which provides corrective feedback for the next generation attempt. This iterative refinement loop continues until syntactically valid specifications and test cases are produced or the error-correction budget is exhausted. Beyond syntactic correctness, we apply soundness and completeness checks that significantly improve compatibility between test cases and specifications (detailed in Section 2.2).
2.1.2. Implementation Generation
@@
An example of a program that successfully passed both phases is shown in Figure 3.
2.2. Soundness and Completeness
@@
Figure 4. Soundness and completeness lemmas constructed from the first test case in Figure 3.
@@
These lemmas evaluate whether the contract is consistent with test cases by checking if it holds pointwise for concrete instantiations. The intuition is that if a specification correctly captures the intended behavior, the contract should be satisfied when instantiated with known correct examples. For each test case, we construct a lemma that asserts the concrete input and output values, then verifies whether the contract holds (Figure 4). If this lemma verifies successfully, it confirms that the contract is consistent with the test case. Conversely, if a soundness lemma fails, it reveals an inconsistency—the contract cannot be satisfied by this known correct example, indicating it is either too restrictive or incorrectly specified.
@@
These lemmas evaluate whether the postconditions uniquely determine outputs for given inputs by attempting to derive a contradiction from alternative outputs. The intuition is that if a specification is sufficiently complete, assuming a different output while maintaining the contract should lead to an inconsistency. For each test case, we construct a lemma that assumes the preconditions, postconditions, and negation of the expected output, then attempts to derive false (Figure 4). If this lemma proves successfully, it confirms no other output can satisfy the postconditions for the given input. However, such proofs often exceed automated theorem provers capabilities without guidance, motivating our perturbation-based alternative described in the next paragraph.
@@
These lemmas provide a practical alternative for evaluating specification completeness by directly testing whether the postconditions accept incorrect outputs. While the contradiction approach attempts to prove that alternative outputs are impossible—a task often too complex for automated provers—the perturbation approach simply checks whether modified outputs satisfy the contract. For each test case, we systematically perturb the output by structurally prompting an LLM while maintaining fixed inputs, then verify whether the contract still holds (Figure 4). If a lemma verifies successfully with a perturbed output, it reveals that the postconditions have insufficient constraints—they accept behavior that should be rejected. Conversely, if the perturbed output fails verification, it suggests the postconditions are appropriately restrictive. This approach trades formal completeness guarantees for practical automated verification.
@@
2.3. Analysis
@@
Figure 5. ATLAS pipeline success rate by TACO-verified difficulty ratings and skill types.
2.3.1. Difficulty Ratings
To evaluate the robustness of our ATLAS pipeline, we analyzed its performance against the difficulty ratings provided by the TACO-verified dataset. The dataset classifies problems into five tiers—EASY, MEDIUM, MEDIUM HARD, HARD, and VERY HARD—enabling us to measure success rate and output characteristics as a function of problem complexity. Results are shown in Figure 5.
@@ 2.3.1. Difficulty Ratings
2.3.2. Skill Set
TACO-verified provides skill set tags for many problems, enabling granular analysis of synthesis capabilities. Among tagged problems, data structure problems achieve the highest success rate at 14.1%, followed by search problems at 13.1%. The pipeline struggles more with sorting and range query problems (6.1% and 4.4%, respectively). Range query problems involve efficiently answering queries about intervals of objects (e.g., determining the maximum sum of a contiguous subarray of fixed length $k$ within a given array), whereas higher-success problem types typically require manipulation of individual elements (e.g., finding the closest-valued node with exactly $k$ descendants). This pattern suggests that formally verifying operations on groups of objects presents a greater challenge than verifying discrete element manipulations, though comprehensive studies would be necessary to substantiate this observation.
3. Supervised Fine-Tuning
3.1. Training
3.1.1. Dataset Curation
Table 1. Multi-task SFT dataset produced by ATLAS.
@@ 3.1.1. Dataset Curation
We hypothesize that fine-tuning for formal verification requires a dataset that teaches more than final answers (cot-reasoning). For agentic applications, the training data must capture the process of problem decomposition, specification, implementation, and correction. We therefore train not only on the final verified Dafny programs ATLAS produces, but also on its execution traces, including intermediate failed attempts. This approach significantly increases the per-sample efficiency of each verified program.
The curated dataset includes: the initial natural language prompt, all intermediate generation attempts for specifications and implementations, LLM-as-a-Judge feedback, and the final verified program. As shown in Table 1, this enables us to extract multiple training tasks from each pipeline run: direct implementation (NL-to-Code Synthesis), specification formalization (NL-to-Spec Synthesis), implementation from specifications (Spec-to-Code Synthesis), assertion synthesis (Proof Infilling), and correction tasks. Correction tasks comprise over 35% of our dataset (Specification Repair and Implementation Repair combined), transforming typically discarded failed attempts into valuable learning opportunities. The Proof Infilling task further leverages successful verifications by teaching the model to reconstruct essential proof annotations such as loop invariants and auxiliary lemmas. These examples are created by stripping proof hints from verified programs. Of the 2.7K verified programs produced by ATLAS, approximately 1K verify automatically without hints, yielding 1.7K training examples.
@@
3.1.2. Execution
This multi-task approach structures our pipeline output into a curriculum for learning formal verification. We fine-tune Qwen 2.5 7B Coder, an open-source LLM for general coding that has limited knowledge of verification-aware languages and syntax. We perform LoRA (hu2021loralowrankadaptationlarge) supervised fine-tuning using the Hugging Face TRL library (trl) on an AWS EC2 instance with 8 NVIDIA H100 GPUs. The model trains for up to 10 epochs using standard cross-entropy loss with gradient clipping. We evaluate on two downstream Dafny benchmarks to assess whether the ATLAS dataset enhances verified code synthesis capabilities.
3.2. Evaluation
3.2.1. DafnyBench
@@
Claude 4.1 Opus [^1] As reported in (vericoding). 89.2
GPT-5111As reported in (vericoding). 72.0
Gemini 2.5 Pro [^1] As reported in (vericoding). 75.3
Claude 3 Opus [^2] As reported in (loughridge2024dafnybenchbenchmarkformalsoftware) at the time of benchmark release. 53.8 65.0 67.8
CodeLlama 7B [^2] As reported in (loughridge2024dafnybenchbenchmarkformalsoftware) at the time of benchmark release. 27.4 28.0 28.0
@@
Figure 6. Evaluation results on DafnyBench.
To assess generalization, we evaluate our fine-tuned model on DafnyBench (loughridge2024dafnybenchbenchmarkformalsoftware), a widely recognized benchmark for Dafny proof hint generation. The benchmark comprises 782 problems requiring assertions and invariants to be filled into ground-truth implementations with specifications.
As shown in Figure 6, ATLAS-enhanced Qwen 2.5 Coder 7B achieves strong single-shot performance at 55.8% Pass@1, surpassing Claude 3 Opus (53.8%), the state-of-the-art at the benchmark’s release (loughridge2024dafnybenchbenchmarkformalsoftware). However, performance plateaus at Pass@5 (56.9%), showing minimal improvement with additional attempts. This contrasts with Claude models, which benefit significantly from multiple attempts—Claude 3 Opus improves from 53.8% to 67.8%. Recent frontier models have established higher performance (vericoding), with Opus 4.1 achieving 89.2% at Pass@10 (vericoding), suggesting our approach may be constrained by the 7B parameter scale of our base model.
Ablation studies reveal that removing specification and completeness checks or test cases has minimal impact on DafnyBench performance, with Pass@10 scores remaining around 56%. This is expected, as DafnyBench primarily focuses on implementing pre-defined contracts rather than open-ended synthesis tasks where these verification components typically prove more valuable (see Section 3.2.2). The benchmark’s nature—having ground-truth specifications and not requiring test case generation—likely undermines the performance boost of soundness and completeness (s&c) checks and test case generation compared to open-ended synthesis tasks. The substantial improvement over the base model Qwen 2.5 Coder 7B (32.4% at Pass@10) and size-matched CodeLlama 7B (28.0% at Pass@10) demonstrates that our ATLAS dataset successfully induces verification-aware capabilities, even if scaling limitations prevent matching larger frontier models.
3.2.2. DafnySynthesis

### v2

@@
Abstract
Large language models have become proficient at generating functional code, but ensuring the output truly matches the programmer’s intent remains difficult. Testing improves trust, yet for safety-critical applications, formal verification provides the only true guarantees through machine-checked proofs. However, verified code remains scarce compared to mainstream languages or mathematical theorem proving, limiting LLM capabilities in this domain. We present ATLAS, an automated pipeline that synthesizes verified programs to address this data bottleneck. Applied to the TACO dataset of Python solutions to LeetCode-style problems, ATLAS generates 2.7K verified Dafny programs, each with high-quality specifications and machine-checked proofs. Through task decomposition, we extract 19K training examples. Fine-tuning Qwen 2.5 7B Coder on this data improves performance from 32.4% to 56.9% on DafnyBench and from 15.8% to 65.8% on DafnySynthesis, demonstrating that synthetic data generation is a viable path to scaling LLM capabilities for formal verification.
Machine Learning, Theorem Proving, Dafny, LLMs, Formal Methods
1 Introduction
Figure 1: An overview of the ATLAS pipeline.
Large language models (LLMs) have transformed software engineering, demonstrating impressive capabilities in code generation, completion, and debugging. On mainstream languages like Python, JavaScript, and Java (Chen et al., 2021; Cassano et al., 2023; Kocetkov et al., 2022), models routinely produce functional, semantically meaningful code that solves the task at hand. Yet in typical workflows, even high-quality generated code comes without guarantees—there is no systematic way to ensure the output truly matches the user’s intent. Unit tests and property-based testing can improve trust, and for most applications this is sufficient.
In safety-critical domains such as cryptographic libraries, medical devices, or aerospace systems, however, "probably correct" is not enough. Formal verification addresses this gap: by requiring precise specifications of intended behavior and mathematically proving implementations satisfy them, verified software eliminates entire classes of bugs by construction. Producing verified code today, however, demands substantial expertise—developers must craft not only the implementation but also specifications and the annotations that guide automated reasoning toward a proof.
Formal verification tools span a spectrum. Interactive theorem provers (ITPs) like Lean and Isabelle require users to construct proofs step-by-step and have seen the most success in mathematical domains. AI-assisted theorem proving in ITPs has advanced rapidly—multiple systems recently achieved gold medal performance on IMO 2025 with formal proofs (Chen et al., 2025; Achim et al., 2025). For software verification, auto-active tools like Dafny (Leino, 2010), F (F* Development Team, 2025), Viper (Müller et al., 2016), and Verus (Lattuada and others, 2023) offer a different tradeoff: users annotate code with specifications and proof hints, and an SMT solver automatically discharges the verification conditions. This automation handles routine reasoning, letting human effort focus on the hard parts. We focus on Dafny and provide further background in Section˜2.
Can similar AI-driven progress be achieved for auto-active verification as has been seen in ITPs? A key obstacle is data scarcity. While code in mainstream languages is abundant, verified code remains rare—Dafny repositories on GitHub are orders of magnitude fewer, and many contain incomplete proofs. The Lean community addressed analogous challenges through synthetic data: autoformalized theorem statements have successfully trained models for theorem proving (Polu et al., 2022; Ying et al., 2025). Inspired by these successes, we tackle the data scarcity problem for Dafny—though our challenge differs: rather than autoformalizing mathematical statements, we must synthesize entire verified programs comprising implementations, specifications, and proof annotations.
To this end, we introduce ATLAS, a pipeline for automatically generating verified programs at scale (Figure˜1). Given a natural language problem description, a reference implementation in a mainstream language (e.g., Python), and test cases, ATLAS synthesizes a corresponding program in a verification-aware language, complete with formal specifications and proof annotations. The pipeline works through iterative refinement, alternating between verifier feedback and LLM-driven repair. To ensure specification quality, we verify that specifications are consistent with test cases and, ideally, strong enough to uniquely determine the expected output for each test input. While the pipeline could be adapted to other input datasets and target languages (e.g., Verus, F) with modest effort, we instantiate it on the TACO-verified dataset (Li et al., 2023)—LeetCode-style problems with Python solutions and unit tests—targeting Dafny.
Figure 2: Evaluation of our fine-tuned Qwen 2.5 7B Coder on DafnyBench and DafnySynthesis, compared with results reported in (Loughridge et al., 2024) and (Misu et al., 2024a) at the time of their release.
[…]
@@ 1 Introduction
  • We present ATLAS, an automated pipeline for synthesizing verified programs from algorithmic problems with reference implementations and test cases.
  • We define soundness and completeness criteria that leverage test cases to ensure specification quality, filtering degenerate solutions.
[…]
  • We show empirically that fine-tuning on our synthesized data yields substantial improvements on DafnyBench and DafnySynthesis.
2 The Dafny Language
Dafny is a verification-aware programming language based on Floyd-Hoare logic (Hoare, 1969). Programs are annotated with specifications in the form of Hoare triples $\{P\}\ s\ \{Q\}$, where $P$ is a precondition (requires), $Q$ is a postcondition (ensures), and $s$ is the program body. Figure 3 shows a simple example.
method Max(a: int, b: int) returns (m: int)
requires true
ensures m >= a && m >= b
ensures m == a $|$ m == b
if a >= b { m := a; } else { m := b; }
Figure 3: A simple Dafny method with specifications.
The verification mechanism relies on weakest precondition calculus. Dafny translates code and specifications into an intermediate representation (Boogie (Barnett et al., 2005)), which computes the weakest precondition wp(s,Q)wp(s,Q)—the necessary and sufficient condition for $s$ to terminate in a state satisfying $Q$. The system then generates a verification condition of the form $P\implies wp(s,Q)$, which is dispatched to the Z3 SMT solver (De Moura and Bjørner, 2008). If Z3 determines the condition is valid, the program is verified.
When Z3 cannot verify a program automatically, users provide ghost code—annotations that guide the solver without affecting runtime behavior. Common forms include assertions (assert P) that establish intermediate facts, loop invariants that characterize state at each iteration, lemmas that encapsulate reusable proof steps, and calculational proofs (calc) that break complex equalities into transitive steps. Unlike interactive theorem provers where users construct explicit proof terms step-by-step, Dafny’s SMT solver handles low-level reasoning automatically while users focus on high-level proof structure.
Dafny has enabled large-scale verification efforts across industry and academia, including AWS’s Cedar authorization engine (Hicks, 2023b), the AWS Cryptographic Material Providers Library (Hicks, 2023a), VMware’s VeriBetrFS file system (Hance et al., 2020), Microsoft’s IronFleet distributed systems (Hawblitzel et al., 2017), and ConsenSys’s Ethereum Virtual Machine specification (ConsenSys, 2023). Notably, AWS rewrote and verified its policy authorization engine in Dafny, which is invoked 1 billion times per second (Chakarov et al., 2025).
3 ATLAS
@@
To address these requirements, we designed ATLAS, a multi-phase pipeline that decouples problem statement generation from verified solution synthesis. The complete process is shown in Figure˜1.
3.1 Overview
@@
Figure 4: A verified Dafny program synthesized by ATLAS, including specification, implementation, and test cases.
We seed the synthesis pipeline with a dataset of programming contest problems paired with solutions in a general-purpose language. This ensures each program has a self-contained implementation and reference solution with input-output test cases. We use TACO-verified (Li et al., 2023), which contains 12.8K programming contest problems with Python reference solutions. We selected TACO-verified for its breadth—covering problems from LeetCode, CodeForces, HackerRank, and others—its unified and rich metadata, and its permissive licensing. Other datasets that could potentially be used for this task include APPS (Hendrycks et al., 2021) and CodeContests (Li et al., 2022), which offer similar problem-solution pairs. However, they may lack the same high-quality metadata and strict test-case quality that TACO-verified provides.
3.1.1 Contract Generation
@@
The Dafny compiler verifies generated specifications and tests for syntax errors. On failure, compiler errors are formatted and passed to an LLM judge, which provides corrective feedback for the next generation attempt. This iterative refinement loop continues until syntactically valid specifications and test cases are produced or the error-correction budget is exhausted. Beyond syntactic correctness, we apply soundness and completeness checks that significantly improve compatibility between test cases and specifications (detailed in Section˜3.2).
3.1.2 Implementation Generation
@@
An example of a program that successfully passed both phases is shown in Figure˜4.
3.2 Soundness and Completeness
@@
Figure 5: Soundness and completeness lemmas constructed from the first test case in Figure˜4.
@@
These lemmas evaluate whether the contract is consistent with test cases by checking if it holds pointwise for concrete instantiations. The intuition is that if a specification correctly captures the intended behavior, the contract should be satisfied when instantiated with known correct examples. For each test case, we construct a lemma that asserts the concrete input and output values, then verifies whether the contract holds (Figure˜5). If this lemma verifies successfully, it confirms that the contract is consistent with the test case. Conversely, if a soundness lemma fails, it reveals an inconsistency—the contract cannot be satisfied by this known correct example, indicating it is either too restrictive or incorrectly specified.
@@
These lemmas evaluate whether the postconditions uniquely determine outputs for given inputs by attempting to derive a contradiction from alternative outputs. The intuition is that if a specification is sufficiently complete, assuming a different output while maintaining the contract should lead to an inconsistency. For each test case, we construct a lemma that assumes the preconditions, postconditions, and negation of the expected output, then attempts to derive false (Figure˜5). If this lemma proves successfully, it confirms no other output can satisfy the postconditions for the given input. However, such proofs often exceed automated theorem provers capabilities without guidance, motivating our perturbation-based alternative described in the next paragraph.
@@
These lemmas provide a practical alternative for evaluating specification completeness by directly testing whether the postconditions accept incorrect outputs. While the contradiction approach attempts to prove that alternative outputs are impossible—a task often too complex for automated provers—the perturbation approach simply checks whether modified outputs satisfy the contract. For each test case, we systematically perturb the output by structurally prompting an LLM while maintaining fixed inputs, then verify whether the contract still holds (Figure˜5). If a lemma verifies successfully with a perturbed output, it reveals that the postconditions have insufficient constraints—they accept behavior that should be rejected. Conversely, if the perturbed output fails verification, it suggests the postconditions are appropriately restrictive. This approach trades formal completeness guarantees for practical automated verification.
@@
3.3 Analysis
Applied to TACO-verified, ATLAS successfully generates 2,751 verified Dafny programs, each with specifications and machine-checked proofs. We analyze pipeline performance across difficulty levels and problem types.
@@ 3.3 Analysis
Figure 6: ATLAS pipeline success rate by TACO-verified difficulty ratings.
Figure 7: ATLAS pipeline success rate by TACO-verified difficulty skill types.
3.3.1 Difficulty Ratings
To evaluate the robustness of our ATLAS pipeline, we analyzed its performance against the difficulty ratings provided by the TACO-verified dataset. The dataset classifies problems into five tiers—EASY, MEDIUM, MEDIUM HARD, HARD, and VERY HARD—enabling us to measure success rate and output characteristics as a function of problem complexity. Results are shown in Figure˜6
@@ 3.3.1 Difficulty Ratings
3.3.2 Skill Set
TACO-verified provides skill set tags for many problems, enabling granular analysis of synthesis capabilities presented in Figure˜7. Among tagged problems, data structure problems achieve the highest success rate at 14.1%, followed by search problems at 13.1%. The pipeline struggles more with sorting and range query problems (6.1% and 4.4%, respectively). Range query problems involve efficiently answering queries about intervals of objects (e.g., determining the maximum sum of a contiguous subarray of fixed length $k$ within a given array), whereas higher-success problem types typically require manipulation of individual elements (e.g., finding the closest-valued node with exactly $k$ descendants). This pattern suggests that formally verifying operations on groups of objects presents a greater challenge than verifying discrete element manipulations, though comprehensive studies would be necessary to substantiate this observation.
4 Supervised Fine-Tuning
The verified programs produced by ATLAS provide a foundation for training models on formal verification. Rather than using only final outputs, we extract multiple training tasks from each pipeline run—including intermediate attempts and repairs—to maximize data efficiency. We describe our dataset curation, training procedure, and evaluation on two Dafny benchmarks.
4.1 Training
4.1.1 Dataset Curation
Table 1: Multi-task SFT dataset produced by ATLAS.
@@ 4.1.1 Dataset Curation
We hypothesize that fine-tuning for formal verification requires a dataset that teaches more than final answers (Wei et al., 2023). For agentic applications, the training data must capture the process of problem decomposition, specification, implementation, and correction. We therefore train not only on the final verified Dafny programs ATLAS produces, but also on its execution traces, including intermediate failed attempts. This approach significantly increases the per-sample efficiency of each verified program.
The curated dataset includes: the initial natural language prompt, all intermediate generation attempts for specifications and implementations, LLM-as-a-Judge feedback, and the final verified program. As shown in Table˜1, this enables us to extract multiple training tasks from each pipeline run: direct implementation (NL-to-Code Synthesis), specification formalization (NL-to-Spec Synthesis), implementation from specifications (Spec-to-Code Synthesis), assertion synthesis (Proof Infilling), and correction tasks. Correction tasks comprise over 35% of our dataset (Specification Repair and Implementation Repair combined), transforming typically discarded failed attempts into valuable learning opportunities. The Proof Infilling task further leverages successful verifications by teaching the model to reconstruct essential proof annotations such as loop invariants and auxiliary lemmas. These examples are created by stripping proof hints from verified programs. Of the 2.7K verified programs produced by ATLAS, approximately 1K verify automatically without hints, yielding 1.7K training examples.
@@
4.1.2 Execution
This multi-task approach structures our pipeline output into a curriculum for learning formal verification. We fine-tune Qwen 2.5 7B Coder, an open-source LLM for general coding that has limited knowledge of verification-aware languages and syntax. We perform LoRA (Hu et al., 2021) supervised fine-tuning using the Hugging Face TRL library (von Werra et al., 2020) on an AWS EC2 instance with 8 NVIDIA H100 GPUs. The model trains for up to 10 epochs using standard cross-entropy loss with gradient clipping. We evaluate on two downstream Dafny benchmarks to assess whether the ATLAS dataset enhances verified code synthesis capabilities.
4.2 Evaluation
4.2.1 DafnyBench
@@
Claude 4.1 Opus [^1] As reported in (Bursuc et al., 2025). 89.2
GPT-5111As reported in (Bursuc et al., 2025). 72.0
Gemini 2.5 Pro [^1] As reported in (Bursuc et al., 2025). 75.3
Claude 3 Opus [^2] As reported in (Loughridge et al., 2024) at the time of benchmark release. 53.8 65.0 67.8
CodeLlama 7B [^2] As reported in (Loughridge et al., 2024) at the time of benchmark release. 27.4 28.0 28.0
@@
Figure 8: Evaluation results on DafnyBench.
To assess generalization, we evaluate our fine-tuned model on DafnyBench (Loughridge et al., 2024), a widely recognized benchmark for Dafny proof hint generation. The benchmark comprises 782 problems requiring assertions and invariants to be filled into ground-truth implementations with specifications.
As shown in Figure˜8, ATLAS-enhanced Qwen 2.5 Coder 7B achieves strong single-shot performance at 55.8% Pass@1, surpassing Claude 3 Opus (53.8%), the state-of-the-art at the benchmark’s release (Loughridge et al., 2024). However, performance plateaus at Pass@5 (56.9%), showing minimal improvement with additional attempts. This contrasts with Claude models, which benefit significantly from multiple attempts—Claude 3 Opus improves from 53.8% to 67.8%. Recent frontier models have established higher performance (Bursuc et al., 2025), with Opus 4.1 achieving 89.2% at Pass@10 (Bursuc et al., 2025), suggesting our approach may be constrained by the 7B parameter scale of our base model.
Ablation studies reveal that removing specification and completeness checks or test cases has minimal impact on DafnyBench performance, with Pass@10 scores remaining around 56%. This is expected, as DafnyBench primarily focuses on implementing pre-defined contracts rather than open-ended synthesis tasks where these verification components typically prove more valuable (see Section˜4.2.2). The benchmark’s nature—having ground-truth specifications and not requiring test case generation—likely undermines the performance boost of soundness and completeness (s&c) checks and test case generation compared to open-ended synthesis tasks. The substantial improvement over the base model Qwen 2.5 Coder 7B (32.4% at Pass@10) and size-matched CodeLlama 7B (28.0% at Pass@10) demonstrates that our ATLAS dataset successfully induces verification-aware capabilities, even if scaling limitations prevent matching larger frontier models.
4.2.2 DafnySynthesis